OpenStudy (anonymous):

write the slope -intercept equation for the line that passes thruough (-7,-5)and has slope-1/7

OpenStudy (anonymous):

plug in the values

OpenStudy (anonymous):

y=mx+c

OpenStudy (anonymous):

y=-x/7-6

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

to find your y-intercept you say -5=(-1/7)*-7+c c=-6 then your slope is m and you just multiply x by m which is x*-1/7